A man has two investments totalling $10,000. On one he receives 8% interest and on the other he receives 4% interest. The 8% investment yields an annual income of $320 more than the 4% investment. How much has he invested at each rate?
Answer by mananth(16946)

A man has two investments totalling $10,000. On one he receives 8% interest and on the other he receives 4% interest. The 8% investment yields an annual income of $320 more than the 4% investment. How much has he invested at each rate?
let one investment be x@8%
and the other be y be @4%
x+y=10000..................(1)
8%x=4%y+320
multiply by 100
8x=4y+32000
8x-4y=32000
/4
2x-y=8000..................(2)
add (1) & (2)
3x=18000
/3
x=6000
$6000@8%
$4000@4%